Thane Company is interested in establishing the relationship between electricity costs and machine hours. Data have been collected and a regression analysis prepared using Excel. The monthly data and the regression output follow:
    

Month Machine Hours Electricity Costs
January 2,600 $ 19,250
February 3,000 $ 22,700
March 2,000 $ 14,350
April 3,200 $ 24,700
May 3,900 $ 29,100
June 3,400 $ 23,700
July 4,200 $ 25,600
August 3,600 $ 23,600
September 2,100 $ 17,200
October 3,800 $ 27,700
November 4,800 $ 32,700
December 4,300 $ 28,600

  

Summary Output
Regression Statistics
Multiple R 0.955
R Square 0.912
Adjusted R2 0.904
Standard Error 1,630.46
Observations 12.00

  

Coefficients Standard Error t Stat P-value Lower 95% Upper 95%
Intercept 4,476.65 1,979.98 2.26 0.05 64.99 8,888.32
Machine Hours 5.76 0.56 10.20 0.00 4.50 7.01

  
If the controller uses the high-low method to estimate costs, the fixed cost portion of the cost equation for electricity costs is: (Round intermediate calculations to 2 decimal places.)

Multiple Choice

  • $1,260.00.

  • $24,640.00.

  • $4,476.65.

  • $1,630.46.
Answer #1

Variable cost = (Cost at highest activity - Cost at lowest activity) /(Highest activity-Lowest activity)

= (32,700-14,350)/(4,800-2,000) = 6.55

Fixed cost = Total cost - Total Variable cost

= 32,700 - (6.55*4800)

= 1260